# Minimax Decision Trees via Martingale Approximations

### v1

@@ A.1 Proof of Theorem 3
A.2 Examples of optimal rates for partition-based martingale approximations
In the following two examples, we show that the geometric rates in Theorem 3 are optimal for the Simons and median martingales.
Example 11 (Rate $r=1/2$ is optimal for the Simons martingale).
$\mathbb{P}\Big{(}U=\sum_{j=1}^{J}(\frac{1-s}{s})^{j}\Big{)}=s^{J+1}(1-s).$
It follows that $U$ is a bounded random variable. Next, we compute that
$\mathbb{E}\Big{[}U\mid U\geqslant\sum_{j=1}^{J}(\frac{1-s}{s})^{j}\Big{]}$ $=\bigg{(}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\bigg{)}^{-1}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\sum_{j=1}^{\ell}(\frac{1-s}{s})^{j}$
$=s^{-(J+1)}\sum_{\ell=J}^{\infty}s^{\ell+1}(2s-1)(1+(\frac{1-s}{s})^{\ell})$
$=\frac{s^{-(J+1)}(2s-1)s^{J+1}}{1-s}+\frac{s^{-(J+1)}(2s-1)s}{1-(1-s)}$
$=\sum_{j=1}^{J+1}(\frac{1-s}{s})^{j}.$
$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=(1-s)^{2k+1}s^{-(k+1)}\asymp(\frac{(1-s)^{2}}{s})^{k}.$
Example 12 (Rate $r=1/2$ is optimal for the median martingale).
Let $s\in(0,1)$ and a bounded random variable $U$ satisfy
Next, we compute that
$=2^{k}\sum_{\ell=k+1}^{\infty}2^{-\ell}(\frac{1}{1-s}-\frac{s^{\ell-1}(2-s)}{1-s})$
$=\frac{1}{1-s}-\frac{2-s}{s(1-s)}\frac{s^{k+1}/2}{1-s/2}$
$=\sum_{j=1}^{k}s^{j-1}.$
Note also that
$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=2^{-k}s^{2k}.$
A.3 Proofs of Theorems 4 and 5
Proof of Theorem 4.
Let $\operatorname{supp}M_{1}=\{x,y\}$ where $x<y$. Take $M>\sup\{x,y\}$. Let $\tau_{M}^{+}$ be the first hitting time to $\{x:x\geqslant M\}$, $\tau_{M}^{-}$ be the first hitting time to $\{x:x\leqslant-M\}$, and $\tau_{M}=\min\{\tau_{M}^{+},\tau_{M}^{-}\}$. It is straightforward to prove (see Zhang et al. (2024)) that for some $r\in(0,1)$ and $C>0$,
Similarly, the same analysis holds for $\tau_{M}^{-}$. Hence, we conclude that $\mathbb{P}(\tau_{M}<\infty)=O(M^{-1}).$ The rest follows line by line as in Zhang et al. (2024), which we sketch for completeness. By Hölder’s inequality, $\mathbb{E}[{\mathbbm{1}}_{\{\tau_{M}<\infty\}}U^{2}]=O(M^{-\delta^{\prime}})$ for some $\delta^{\prime}>0$. With the choices $M=r^{-k/(2+\delta^{\prime})}$ and $q=r^{\delta^{\prime}/(2+\delta^{\prime})}\in(0,1)$, we have that for some $C>0$,
Proof of Theorem 5.
To see this, without loss of generality we start from $U$ supported on $[0,1]$ (meaning that $\mathbb{P}(U\in[0,1])=1$) whose density takes values in $[1,1+\varepsilon]$ (we may normalize afterwards to a probability measure, which we omit). Then for any interval $[a,b]\subseteq[0,1]$,
Therefore,
$\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(b-a)^{3}.$
Similarly,
It follows that the split point $x_{*}\in[0,1]$ must satisfy
Hence, as $\varepsilon\to 0$, the ratio becomes
as desired. ∎
A.4 Proofs of results from Section 3
Lemma A.1.
Let $\delta>0$. Then for any coupling $(X,Y)$ with $X\geqslant\mathbb{E}[Y]$, it holds that
@@ A.4 Proofs of results from Section 3
If $\sup\operatorname{supp}X/\inf\operatorname{supp}X\geqslant(1-\sqrt{(1+\delta)/(C\delta)})^{-1}$, the second term is always non-negative and hence (39) holds. Otherwise, since $C=(1+\delta)^{3}/\delta^{3}$,
$(1+\delta)(\inf\operatorname{supp}X)^{2}\geqslant(1+\delta)(1-\sqrt{(1+\delta)/(C\delta)})(\sup\operatorname{supp}X)^{2}\geqslant(\sup\operatorname{supp}X)^{2}.$
@@
It remains to establish (40). Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ from the cyclic minimax construction. Note that by construction, #$\#J_{k}\leqslant 2^{k}$ and $M_{k}\mid\mathbf{X}\in I_{j}$ is a piecewise constant random variable, which takes a constant value $y_{k,j}$ inside the interval $I_{j}$, so $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$. By definition, $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$. Define $g_{j}=g|_{I_{j}}$, so that $\left\lVert g\right\rVert_{\mathrm{TV}}=\sum_{j}\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$. Define $\mathrm{Ran}(g_{j})=[\inf g_{j},\sup g_{j}]$ and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
The key step is to decompose the local mean-squared error $\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ into two parts for each $j$, with the help of Lemma A.1. Next, we claim that
@@
    $y_{k,j}\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and $V_{j}=(1+\delta^{-1})\,\mathbb{E}[(g(\mathbf{X})-y_{k,j})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$.
@@
    $y_{k,j}\not\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and $V_{j}=p_{j}(\Delta g_{j})^{2}(1+\delta)^{3}/\delta^{3}$.
In case (i), (43) follows immediately from the polarization identity $(a+b)^{2}\leqslant(1+\delta)a^{2}+(1+\delta^{-1})b^{2}$. On the other hand, for case (ii), (43) follows by applying Lemma A.1 to the coupling $(g(\mathbf{X}),Y)\mid\mathbf{X}\in I_{j}$ and using that $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$. This completes the proof of (43).
@@
To control $\sum_{j}V_{j}$, we apply the same Hölder’s inequality argument as in the proof of Theorem 3 in the minimax case. We have
@@
To further bound the right-hand side of (46), we recall from (21) that as a consequence of the marginally atomless property,
@@
This is immediate for case (ii) above (when $y_{k,j}\not\in\mathrm{Ran}(g_{j})$). By definition and using $y_{k,j}\in\mathrm{Ran}(g_{j})$, we have for case (i) above (when $y_{k,j}\in\mathrm{Ran}(g_{j})$),
@@
[…]
@@
The bound (41) can be improved to
@@
The only difference from the proof of Theorem 6 is (47), since for non-marginally atomless measures (20) and (21) may not hold. Instead, by our assumptions, (20) is be replaced by
$\begin{split}&\hskip 14.22636pt\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}\}}],\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}\}}]\Big{\}}\\ &\leqslant\frac{1}{2}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]+\frac{M^{2}}{N},\end{split}$
because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $M^{2}/N$. As a consequence, if we denote by
$u_{k}:=\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}],$
@@
$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{M^{2}}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=\mathrm{Var}(Y).$
To solve this, let $v_{k}:=u_{k}-2M^{2}/N$. Suppose that $u_{0}\geqslant 2M^{2}/N$. Then $v_{k}$ satisfies $v_{k+1}\leqslant v_{k}/2$ and $v_{0}\leqslant\mathrm{Var}(Y)$. It follows that $v_{k}\leqslant 2^{-k}\mathrm{Var}(Y)$ for all $k\geqslant 0$, and hence
$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]=u_{k}\leqslant 2^{-k}\mathrm{Var}(Y)+\frac{2M^{2}}{N}.$ (51)
It is also easy to check that (51) is also satisfied for the case $u_{0}<2M^{2}/N$, that is, (51) holds in general. Replacing (47) by (51) leads to
$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg{(}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\Big{(}\left\lVert g\right\rVert_{\mathrm{TV}}(\mathrm{Var}(Y)+2^{k+1}\frac{M^{2}}{N})\Big{)}^{2/3}\bigg{)}.$
@@
$\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\Big{(}2^{k+1}\frac{M^{2}}{N}\Big{)}^{2/3}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{2}}{3}+\frac{2^{k+2}}{3}\frac{M^{2}}{N},$
@@
Note that if $\mathbf{X}_{*}$ is marginally atomless, the assumption (27) for the coupling $(\mathbf{X},Y)$ in Theorem 9 holds $\mathbb{P}_{*}$-almost surely. Consequently, the proof is almost verbatim compared to Theorem 4.3 of Klusowski and Tian (2024), by applying our Theorem 9 instead of Theorem 4.2 therein. The only difference here is the extra term
$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}$
appearing in (28), where we remind the reader that $M=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$. Nevertheless, since the proof of Theorem 4.3 of Klusowski and Tian (2024) deals first with the case with bounded data, we may simply replace $M$ by that bound. Specifically, the first part of the proof of Theorem 4.3 of Klusowski and Tian (2024) assumes $\max_{i}|Y_{i}|\leqslant U$. We may set $M=2U$ and add the term
$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}\leqslant\frac{2^{k+5}U^{2}}{3N}$

### v2

@@ A.4 Proof of Theorem 16
Appendix B Proofs of main results
B.1 Proofs of results in Section 2
Proof of Lemma 2.1.
We take arbitrary $a<b$ and define events $E=E_{1}\cup E_{2}$, where $E_{1}=\{\mathbf{X}\in A,\,X_{j}<a\}$ and $E_{2}=E\setminus E_{1}=\{\mathbf{X}\in A,\,a\leqslant X_{j}<b\}.$ By the total variance formula,
Rearranging gives
and therefore $\varphi_{L}$ is non-decreasing; and similarly $\varphi_{R}$ is non-increasing. The continuity of $\varphi_{L}$ and $\varphi_{R}$ follows from the marginally atomless condition. ∎
We also need the following risk decomposition lemma.
Lemma B.1.
Let $\delta>0$. Then for any joint distribution $(X,Y)$ with $X\geqslant\mathbb{E}[Y]$, it holds that
@@ B.1 Proofs of results in Section 2
If $\inf\operatorname{supp}X=0$ or $\sup\operatorname{supp}X/\inf\operatorname{supp}X\geqslant(1-\sqrt{(1+\delta)/(C\delta)})^{-1}$, the second term is always non-negative and hence (44) holds. Otherwise, since $C=(1+\delta)^{3}/\delta^{3}$,
$(1+\delta)(\inf\operatorname{supp}X)^{2}\geqslant(1+\delta)(1-\sqrt{(1+\delta)/(C\delta)})(\sup\operatorname{supp}X)^{2}=(\sup\operatorname{supp}X)^{2}.$
@@
It remains to establish (45). Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ by the cyclic minimax construction. Note that by construction, $J_{k}\leqslant 2^{k}$ and $M_{k}\mid\mathbf{X}\in I_{j}$ is a piecewise constant random variable, which takes a constant value $y_{k,j}$ inside the box $I_{j}$, so $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$. By definition, $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$. Define $g_{j}=g|_{I_{j}}$, so that $\left\lVert g\right\rVert_{\mathrm{TV}}=\sum_{j}\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$. Define $\mathrm{Ran}(g_{j})=[\inf g_{j},\sup g_{j}]$ and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
The key step is to decompose the local mean squared error $\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ into two parts for each $j$, with the help of Lemma B.1. Next, we claim that
@@
    $y_{k,j}\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
    $V_{j}=\min\big{\{}(1+\delta^{-1})\,\mathbb{E}[(g(\mathbf{X})-y_{k,j})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}],\,\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\big{\}}.$
@@
    $y_{k,j}\not\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
    $V_{j}=\min\Big{\{}\frac{(1+\delta)^{3}}{\delta^{3}}\,p_{j}(\Delta g_{j})^{2},\,\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\Big{\}}.$
In case (i), (48) follows immediately from the polarization identity $(a+b)^{2}\leqslant(1+\delta)a^{2}+(1+\delta^{-1})b^{2}$. On the other hand, for case (ii), (48) follows by applying Lemma B.1 to the joint distribution $(g(\mathbf{X}),Y)\mid\mathbf{X}\in I_{j}$ and using that $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$. This completes the proof of (48).
@@
To control $\sum_{j}V_{j}$, we apply Hölder’s inequality. We have
@@
To further bound the right-hand side of (51), we first recall from (11) that as a consequence of the marginally atomless property,
@@
This is immediate for case (ii) above (when $y_{k,j}\not\in\mathrm{Ran}(g_{j})$). By definition, we have for case (i) above (when $y_{k,j}\in\mathrm{Ran}(g_{j})$),
@@
where $\lfloor\cdot\rfloor$ is the floor function. Combined with (53), we arrive at
@@
In the proof of Theorem 2, the bound (46) can be improved to
@@
Recall that $\Delta Y:=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$. The only difference from the proof of Theorem 2 is (52), since for non-marginally atomless measures (10) and (11) may not hold. Instead, by our assumptions, (10) is be replaced by
$\max\Big{\{}\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big{\}}\leqslant\frac{1}{2}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]+\frac{(\Delta Y)^{2}}{N},$
because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $(\Delta Y)^{2}/N$. As a consequence, if we denote by
$u_{k}:=\max_{A\in\pi_{k}}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}],$
@@
$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{(\Delta Y)^{2}}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=\mathrm{Var}(Y).$
To solve this, let $v_{k}:=u_{k}-2(\Delta Y)^{2}/N$. Suppose that $u_{0}\geqslant 2(\Delta Y)^{2}/N$. Then $v_{k}$ satisfies $v_{k+1}\leqslant v_{k}/2$ and $v_{0}\leqslant\mathrm{Var}(Y)$. It follows that $v_{k}\leqslant 2^{-k}\mathrm{Var}(Y)$ for all $k\geqslant 0$, and hence the maximum risk among the nodes at level $k$ is
$\max_{A\in\pi_{k}}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]=u_{k}\leqslant 2^{-k}\mathrm{Var}(Y)+\frac{2(\Delta Y)^{2}}{N}.$ (56)
It is also easy to check that (56) is also satisfied for the case $u_{0}<2(\Delta Y)^{2}/N$, that is, (56) holds in general. Replacing (52) by (56) leads to
$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg{(}(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\Big{(}\left\lVert g\right\rVert_{\mathrm{TV}}(\mathrm{Var}(Y)+2^{k+1}\frac{(\Delta Y)^{2}}{N})\Big{)}^{2/3}\bigg{)}.$
@@
$\left\lVert g\right\rVert_{\mathrm{TV}}^{2/3}\Big{(}2^{k+1}\frac{(\Delta Y)^{2}}{N}\Big{)}^{2/3}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}^{2}}{3}+\frac{2^{k+2}}{3}\frac{(\Delta Y)^{2}}{N},$
@@
Note that if $\mathbf{X}_{*}$ is marginally atomless, the assumption (22) for the joint distribution $(\mathbf{X},Y)$ in Theorem 6 holds $\mathbb{P}_{*}$-almost surely. Consequently, the proof is almost verbatim compared to Theorem 4.3 of Klusowski and Tian (2024), by applying our Theorem 6 instead of Theorem 4.2 therein. The only difference here is the extra term
$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}(\Delta Y)^{2}}{3N}$
appearing in (23), where we remind the reader that $\Delta Y=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$. Nevertheless, since the proof of Theorem 4.3 of Klusowski and Tian (2024) deals first with the case with bounded data, we may simply replace $\Delta Y$ by that bound. Specifically, the first part of the proof of Theorem 4.3 of Klusowski and Tian (2024) assumes $\max_{i}|Y_{i}|\leqslant U$. We may set $\Delta Y=2U$ and add the term
$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}(\Delta Y)^{2}}{3N}\leqslant\frac{2^{k+5}U^{2}}{3N}$